OpenStudy (anonymous):

What place in this world can have their temperatures Fahrenheit and Celsius equal?

OpenStudy (anonymous):

F = 1.8C + 32 x = 1.8x + 32 x - 1.8x = 1.8x + 32 - 1.8x -0.8x = 32 -8x = 320 -x = 40 x = -40 The only place that might be -40 degrees cold must be Anartica

OpenStudy (anonymous):

Oh kayyy, cool